OpenStudy (anonymous):

solve for the absolute value |3x+9|-12=-8

OpenStudy (anonymous):

|3x+9|-12=-8 |3x+9|=-8+12 |3x+9|=4 --------- This is the hard part where you will have two answers. +(3x+9)=4 vs -(3x+9)=4 Starting with +(3x+9)=4 +(3x+9)=4 3x+9=4 3x=4-9 3x=-5 x=-5/3 Second Equation with -(3x+9)=4 -(3x+9)=4 -3x-9=4 -3x=4+9 -3x=13 x=13/-3 So your answers are -5/3 and -13/3

OpenStudy (anonymous):

I worked this question many times, but no matter what i did, i got the same answer you got. But what's bugging me is that my lecturer said that when solving absolute value with inequalities,of the two answers, one must reflect a positive value and the other must reflect a nagative value. Babiimuffin, i thank you for the energy you dedicated towards helping me to get a solution to my question.

OpenStudy (anonymous):

your instructor is partially correct the reason both your answers are negative is because in one of them your subtraction of numbers makes the number negative so one would be positive it wasn't that 3x=4-9 making the right side negative